OpenStudy (anonymous):

the value of \[1^{125}+1^{^{125}}\]

OpenStudy (anonymous):

2

OpenStudy (lalaly):

1 raised to any number is 1

OpenStudy (anonymous):

why ? please give an explanation?

OpenStudy (lalaly):

1^2 = 1 x1 = 1 1^3 = 1x1x1 = 1 1^4 = 1x1x1x1 =1 1^5 =1x1x1x1x1 =1 and so on

OpenStudy (king):

so answer is 2!

OpenStudy (anonymous):

why ? give an explanation king?

OpenStudy (anonymous):

ok..tanks lalaly>?

OpenStudy (lalaly):

are u kidding? 1^125 = 1 so 1+1 =2

OpenStudy (anonymous):

i'm not kidding.. i don't kno this question

OpenStudy (lalaly):

i told u, 1^(anything) = 1 so u are then left with 1+1

OpenStudy (anonymous):

ok...i agree

OpenStudy (lalaly):

so 1+1=?

OpenStudy (anonymous):

2

OpenStudy (lalaly):

\[1^{125}=1\] \[1^{125} + 1^{125} = 1+1 =2\]

OpenStudy (anonymous):

ok..i agree thanks lalaly?

OpenStudy (lalaly):

you are welcome yecci:)

OpenStudy (anonymous):

:)